C = 450 + 0.4Y
I = 350
G = 150
X = 70
Z = 35 + 0.1Y
T = 0.15Y
Yf = 1550
1
Expert's answer
2020-11-30T16:38:42-0500

Y=C+I+G+X-Z

Y=450+0.4(1550-0.15Y)+350+150+70-(35+0.1Y)

Y=1605-0.16Y

Y=1383.62